# Quantum cryptography beyond key distribution: theory and experiment

### v2

@@ V.1 Special relativity and trusted agents
For completeness, we note the existence of other mistrustful protocols basing their security on spacetime constraints, such as quantum oblivious transfer Pitalúa-García (2016, 2019), device-independent bit commitment Adlam and Kent (2015), multiparty coin flipping Pitalúa-García (2021) and zero-knowledge proofs in the multi-prover setting Ben-Or et al. (1988); Chailloux and Leverrier (2017); Grilo et al. (2019); Alikhani et al. (2021).
@@ VI.1 Public-key quantum money
Problem $(i)$ was already considered in Bennett et al. (1983), where the quantum token |$\ket{\psi_{s}}$ depends on a public serial number $s$. Instead of storing a random sequence of quantum states for each serial number, the authors proposed the use of a quantum-secure, pseudorandom function $f_{k}$, where the key $k$ is only known to the bank and the output $f_{k}(s)$ determines the states to be prepared and the information required for verification. From a security perspective, an adversary with unlimited computational power can recover $k$ and break the scheme, but any computationally-bounded adversary cannot, since this would imply breaking the pseudorandomness property of $f_{k}(s)$.
Problems $(ii)$ and $(iii)$ may be addressed by designing a protocol in which all parties, not only the bank, are able to verify the quantum money. First introduced in Aaronson (2009); Farhi et al. (2012); Aaronson and Christiano (2012), this concept is known as public-key quantum money. The basic idea is to produce a state |$\ket{\psi}$, such that the bank can publish a description of a quantum circuit that lets other parties recognize |$\ket{\psi}$, but does not allow them to prepare or even copy |$\ket{\psi}$. This may not be achieved in an i.t.-secure way, but becomes possible with computational assumptions. Whenever this is the case, it is crucial to find assumptions that we are confident cannot be broken (such as the minimal assumptions introduced in Section II.6). A lot of research effort is dedicated to weakening assumptions and eventually resorting to so-called “standard cryptographic assumptions”, such as the existence of a quantum random oracle, the hardness of Learning With Errors (LWE) or the existence of post-quantum secure one-way functions. For public-key quantum money, a scheme whose security relies on the hardness of a knot theory problem – a non-standard assumption –, was provided in Farhi et al. (2012), and its security is yet to be proven (or disproven). In other works, the quantum money state takes the form
|$\ket{A}:=\frac{1}{\sqrt{|A|}}\sum_{x\in A}\ket{x},$ (12)
i.e. an equal superposition of all bit strings of a subset $A$ that consists of half the possible bit strings Aaronson and Christiano (2012). The classical description of the subset $A$ remains hidden from the users and is only known to the bank. The bank also publicizes the details of how to check that the note is correct using calls to membership oracles, with essentially the property that $V_{A}\ket{A}=\ket{A}$. While verification requires a single call to the verification oracle $V_{A}$, an adversary could cheat with an exponential number of queries to that oracle. Note that giving access to a program without leaking information about its inner workings can be achieved with a cryptographic notion called indistinguishability obfuscation (IO).
Another major development appeared in Zhandry (2021), where a stronger notion of public-key quantum money became possible, subsequently termed “public-key semi-quantum money”. In addition to enabling public verification, the bank becomes fully classical, simply instructing the untrusted participants to generate the quantum money themselves. Based on the idea of quantum lightning Zhandry (2021), the trick is to prepare an equal superposition of inputs as shown in Section II.5, using a collision-resistant hash function $H$. Specifically, one can generate a state |$\ket{\psi_{y}}=\sum_{x|H(x)=y}\ket{x}\ket{y}$ for a random $y$. It is then easy to see that nobody can prepare the same state twice without breaking the collision resistance of $H$. For its use in quantum money, the crucial property is to be able to verify it, which is possible if one can distinguish between the states |$\ket{x}$ and |$\ket{\psi_{y}}$, a property known as non-collapsible collision-resistant hash functions [^3] It is an open question whether such functions exist.. A construction for public-key semi-quantum money based on weaker assumptions, namely IO and LWE, was later developed Shmueli (2022).
@@ VI.2 Tokenized and one-shot signatures
Quantum tokens as digital signatures were introduced in Ben-David and Sattath (2023). To get a little intuition, consider the simple case of signing a one-bit message. Alice selects a subspace $A$ and prepares the corresponding quantum token |$\ket{A}$ exactly as in Eq. (12) from Section VI.1. Bob can sign the message $m=0$ by returning an $x$ that belongs to the subspace $A$, which he obtains by measuring the quantum token in the computational basis. To sign the message $m=1$, Bob must return an $x$ that does not belong to $A$, which he obtains by first applying the Hadamard $H$ to all qubits to obtain |$\ket{A^{\bot}}$ and then measuring the register in the computational basis. Bob then returns the pair $(m,x)$ as a signature and Alice checks if $x$ is compatible with the message. To turn this into a publicly verifiable signature scheme, Alice can use the same methods as in Section VI.1, and give obfuscated access to the membership test, i.e. can allow other parties to check if a given $x$ belongs to $A$ or $A^{\bot}$ without revealing the sets.
One-shot signatures were introduced in Amos et al. (2020). In this setting, we require that even Alice, the signing authority, cannot sign a second message corresponding to the same public key. Interestingly, by achieving this, one can now delegate the generation of the quantum token to Bob, making Alice completely classical. Alice picks a collision-resistant hash function $H$ and instructs Bob to generate an equal superposition of a random image using the trick of Section II.5. The first register |$\ket{\psi_{y}}=\sum_{x|H(x)=y}\ket{x}$ is the token. To sign a message, Bob needs to output a pre-image such that the first bit of the string is equal to the message he wants to sign. To construct an algorithm that can “select” a pre-image matching the message that Bob wants to sign without any secret information, the hash function requires extra properties, such as being non-collapsible. The triplet $(m,x,y)$ materializes the signature, and anyone can verify its validity by checking the first bit of $x$ and computing $H(x)$ to check that it is equal to $y$.
@@ VI.3 Public-key cryptography
When quantum keys are involved, we need an extra generation procedure $\operatorname{Genp}$ to produce public keys |$\ket{pk}$ from secret keys $sk$. It is generally assumed that |$\ket{pk}$ is a pure quantum state. The encryption process $\operatorname{Enc}$ maps the public key |$\ket{pk}$ and classical message $m$ to a ciphertext $\rho_{c}$ (generally a mixed state) and $\operatorname{Dec}$ maps the private key $sk$ and ciphertext $\rho_{c}$ back to the classical message $m$.
An example scheme from Barooti et al. (2023) assumes the existence of a quantum-safe pseudorandom function (PRF), which can be obtained from an arbitrary quantum-safe one-way function Zhandry (2012). Given a PRF ensemble $\{f_{k}\}_{k}$, the public key consists of a pair of pure quantum states (|$(\ket{qpk_{0}},\ket{qpk_{1}})$ and the secret key consists of a pair of strings $(dk_{0},dk_{1})$ such that, for all $b\in\{0,1\}$,
|$\ket{qpk_{b}}=\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}\ket{x,f_{dk_{b}}(x)}.$ (13)
To encrypt a bit $b$, one measures |$\ket{qpk_{b}}$ in the computational basis. The result is $(x,f_{dk_{b}}(x))$ for some uniformly random $x\in\{0,1\}^{n}$ and this is the ciphertext, i.e., $(qc_{0},qc_{1})=(x,f_{dk_{b}}(x))$.
@@ VII Multipartite quantum cryptography
In the context of quantum networks, most applications will involve more than two parties interacting, which justifies a section dedicated to multipartite quantum primitives. Naturally, the protocols we study here either extend two-party quantum primitives to the multipartite setting Pitalúa-García (2021); Ambainis et al. (2004), or intrinsically require more than two parties, such as secret sharing, data hiding, Byzantine agreement, randomized leader election and electronic voting. Since QKD and its multipartite extension are beyond the scope of this review, we omit the study of quantum conference key agreement and refer the interested Reader to the review from Murta et al. (2020).
@@
|$\ket{\psi^{(n)}_{shared}}=\frac{1}{\sqrt{2}}\left(\ket{0}^{\otimes n}+\ket{1}^{\otimes n}\right).$ (14)
@@ VII.3 Byzantine agreement
The Byzantine agreement problem has widespread applications, especially in distributed computation where the processors involved might be faulty or dishonest. Initially formulated by Lamport et al. (1982), it may be stated as follows: several divisions of the Byzantine army, each led by an independent general, are camped around an enemy city. Conditioned on the enemy’s actions, all Byzantine generals should agree on a common plan of action (eg: attack or retreat), even in the presence of potential traitors attempting to scramble the plan. Given the geometry of the situation, generals can only communicate via messenger. In the case of one commanding general sending an order to their $(n-1)$ lieutenant generals, the following conditions must be met for the action plan to succeed Lamport et al. (1982):
@@
A cheat-sensitive version of Byzantine agreement was proposed in the $3$-general setting by adding quantum communication to the pairwise authenticated classical channels Fitzi et al. (2001). In the honest protocol, the commanding general distributes $j$ copies of a $3$-partite entangled state |$\ket{\psi}$ consisting of qutrits spanned by {|$\{\ket{0},\ket{1},\ket{2}\}$:
@@ VIII.1 Private channels
An alternative encryption scheme uses classical keys and produces a quantum ciphertext Ambainis et al. (2000); Boykin and Roychowdhury (2003). If |$\ket{\psi}$ is a qubit to be encrypted, then the encryption scheme uses a pair of random bits $(a,b)$ as the classical key and |$\ket{\psi_{ciph}}=X^{a}Z^{b}\ket{\psi}$ as the quantum ciphertext. For any party who is completely ignorant about the key, the security guarantee is that the mixed state associated with |$\ket{\psi_{ciph}}$ is indeed the completely mixed state, i.e. appears random to that party. Since the latter is independent of the input |$\ket{\psi}$, this scheme is also perfectly secret in the sense of Shannon, and known as the quantum Vernam cipher.
@@ VIII.2 Authentication
In a quantum setting, these notions have to be adapted because the sender does not have several copies of |$\ket{m}$ and thus it cannot both be used to create |$\ket{t}$ and sent alongside with it. In this case, we can think of the authentication process as a quantum mapping of a message |$\ket{m}$ and key $k$ to its authenticated form $A_{k}\ket{m}$, for some isometry $A_{k}$. The receiver will use a measurement dependent on $k$ to determine the validity of the message and recover the message itself.
@@
The approach proposed in Barnum et al. (2002) is to encrypt the message |$\ket{m}$ as in Section VIII.1, mix it with some fixed maximally-mixed check qubits and apply a quantum error-correcting code. The quantum encryption of the message makes it indistinguishable from the check positions. The check positions are present to detect any significant modification of the transmission whereas the quantum error-correcting code is present to recover from any minor modification of the message.
@@ VIII.3 Anonymous communication
Figure 6: Anonymous multipartite entanglement generation. Initially proposed as a subroutine for quantum conference key agreement and taken from Hahn et al. (2020), this diagram pictures a sender $A$ distributing an $n$-partite GHZ state to herself and $(n-1)$ other parties. To the left of arrow $(1)$, $A$ uses a classical protocol to anonymously notify a subset of $m<n$ blue receivers $\{B_{1},B_{2},...,B_{m}\}$ that she will later establish a secret key with them Broadbent and Tapp (2007). To the right of arrow $(1)$, all parties are colored in pink since the shared $n$-partite GHZ state is agnostic of the receivers’ identities. To reach the right of arrow $(2)$, all unnotified parties measure their subsystem in the $X$ basis and broadcast their measurement outcomes in a random order (or through a simultaneous broadcast channel), while the notified parties declare random bits to preserve their anonymity. $A$ applies a phase flip to her subsystem if the parity of the unnotified parties’ outcomes is odd, yielding the desired $(m+1)$-partite GHZ state.
@@ IX.1 Blind computation
The basic trick to make the protocol blind is the following observation. Within MBQC, the effect of measuring a single qubit in the basis |±$\ket{\pm_{\phi}}$ on the standard resource state – where qubits are initialized in the $\ket{+}$ state and then entangled with $\wedge Z$ gates, is identical to measuring a resource state (that is similarly rotated) in the basis |±$\ket{\pm_{(\phi+\theta)}}$ – where the qubits are prepared in the |$\ket{+_{\theta}}$ state prior to getting entangled. The client, by deciding and hiding the pre-rotation angle $\theta$, can instruct the server to measure in the |±$\ket{\pm_{(\phi+\theta)}}$ basis, without revealing any information on the angle $\phi$ – essentially $\theta$ “one-time-pads” the angle $\phi$. However, as stated above, the angles $\phi$ fully determine the unitary performed, and thus the server is completely blind as to the computation they are performing. The actual instructions have the client asking the server to measure in an angle with an extra $r\pi$ flip, where $r$ is a random bit that is secret to the server. The purpose of this bit is to hide (one-time-pad) the measurement outcomes.
@@
Figure 7: Universal blind quantum computing. First proposed in Broadbent et al. (2009), this figure was taken from Gheorghiu et al. (2019), and reproduced with permission from the authors. For each position $(i,j)$ in the rectangular cluster state, the client prepares a random qubit of the form |$\ket{+_{\theta}}_{ij}$ as defined in Eq. (17), sends it to the server, and requires them to prepare the rotated graph state. The client then instructs the server to perform a measurement $M(\delta_{ij})$ in a basis with angle $\delta_{ij}$, which cancels the pre-rotation $\theta_{ij}$ and performs the desired computation dictated by $\phi_{ij}$ in a blind way. Note that the extra term $r_{ij}\pi$ essentially one-time pads the measurement outcome $b_{ij}$.
@@ IX.2 Verifiable computation
In Fitzsimons and Kashefi (2017), the traps were isolated single qubits of the form |$\ket{+_{\theta}}$, where the isolation was achieved by surrounding trap qubits with “dummy” qubits prepared in the computational basis {|$\{\ket{0},\ket{1}\}$. The trap qubits are not entangled with the rest of the resource state because their dummy neighbours are in the computational basis and the $\wedge Z$ gates do not entangle them. Therefore, measuring trap qubits in the suitable basis returns a deterministic result. To hide the positions of the traps, and to ensure that the probability of accepting a corrupted computation becomes negligible, quantum error correcting codes and a graph that scales quadratically compared to the non-verifiable version are used.
@@ IX.4 Fully homomorphic encryption
Our task is to construct a quantum version of this scenario, where, instead of evaluating a function on a classical message, we wish to apply a unitary on a quantum state: the client sends $\textrm{Enc}\ket{\psi}$ to the server and expects to receive $\textrm{Enc}(U\ket{\psi})$ back. Such a task was achieved for unitaries with a constant (low) $T$-gate depth in Broadbent and Jeffery (2015). Since this construction does not handle arbitrary unitaries however, it is not a “fully” homomorphic scheme. One first observation is that, for Clifford gates, homomorphic evaluation can be easily performed by applying the gates and modifying the “corrections”, i.e. modifying the one-time-pad on the quantum state. In Broadbent and Jeffery (2015), the authors construct a gadget in which the client hands an (unknown) quantum state to the server, enabling the server to apply a $T$-gate on the encrypted quantum state. While this can be done non-interactively, the cost of the overall method scales exponentially with the $T$-depth of the unitary (using a classical FHE scheme in parallel).
@@
On the experimental side, a proof-of-principle computation on encrypted quantum data was performed using polarization-encoded single photons and linear optics Fisher et al. (2014). In this scheme, the client is only required to prepare and send single qubits, encrypted through the quantum equivalent of a one-time pad (Section VIII.1). Since client and server agree on the computation beforehand, the overhead in terms of auxiliary qubits is proportional to the number of non-Clifford gates in the computation. Single-qubit Clifford gates are naturally performed using a set of waveplates, while the CNOTs require two-photon interference at a partially polarizing beamsplitter (giving a success probability of $1/9$) and the non-Clifford gates are actively implemented using various Pockels cells to prepare the auxiliary qubit. While a full practical security analysis is not provided, the leakage of information due to imperfect quantum encryption on the client’s side and multiphoton emission from the SPDC source is quantified. Two years later, a similar protocol was implemented in the continuous-variable setting at telecom wavelengths, showing that security can be achieved with up to $10$km of lossy channel between the client and the server Marshall et al. (2016).

### v3

@@ V.1 Special relativity and trusted agents
For completeness, we note the existence of other mistrustful protocols basing their security on spacetime constraints, such as quantum oblivious transfer Pitalúa-García (2016, 2019), device-independent bit commitment Adlam and Kent (2015), multiparty coin flipping Pitalúa-Garc´ıa (2021) and zero-knowledge proofs in the multi-prover setting Ben-Or et al. (1988); Chailloux and Leverrier (2017); Grilo et al. (2019); Alikhani et al. (2021).
@@ VI.1 Public-key quantum money
[…]
i.e. an equal superposition of all bit strings of a subset $A$ that consists of half the possible bit strings Aaronson and Christiano (2012). The classical description of the subset $A$ remains hidden from the users and is only known to the bank. The bank also publicizes the details of how to check that the note is correct using calls to membership oracles, with essentially the property that $|AV_{A}\ket$$\ket{A}$. While verification requires a single call to the verification oracle $V_{A}$, an adversary could cheat with an exponential number of queries to that oracle. Note that giving access to a program without leaking information about its inner workings can be achieved with a cryptographic notion called indistinguishability obfuscation (IO).
[…]
@@ VI.2 Tokenized and one-shot signatures
[…]
@@ VI.3 Public-key cryptography
[…]
@@ VII Multipartite quantum cryptography
In the context of quantum networks, most applications will involve more than two parties interacting, which justifies a section dedicated to multipartite quantum primitives. Naturally, the protocols we study here either extend two-party quantum primitives to the multipartite setting Pitalúa-Garc´ıa (2021); Ambainis et al. (2004), or intrinsically require more than two parties, such as secret sharing, data hiding, Byzantine agreement, randomized leader election and electronic voting. Since QKD and its multipartite extension are beyond the scope of this review, we omit the study of quantum conference key agreement and refer the interested Reader to the review from Murta et al. (2020).
@@
[…]
@@ VII.3 Byzantine agreement
The Byzantine agreement problem has widespread applications, especially in distributed computation where the processors involved might be faulty or dishonest. Initially formulated by Lamport et al. (1982), it may be stated as follows: several divisions of the Byzantine army, each led by an independent general, are camped around an enemy city. Conditioned on the enemy’s actions, all Byzantine generals should agree on a common plan of action (eg: attack or retreat), even in the presence of potential traitors attempting to scramble the plan. Given the geometry of the situation, generals can only communicate via messenger. In the case of one commanding general sending an order to their (n1)(n-1) lieutenant generals, the following conditions must be met for the action plan to succeed Lamport et al. (1982):
@@
[…]
@@ VIII.1 Private channels
[…]
@@ VIII.2 Authentication
[…]
@@
[…]
@@ VIII.3 Anonymous communication
Figure 6: Anonymous multipartite entanglement generation. Initially proposed as a subroutine for quantum conference key agreement and taken from Hahn et al. (2020), this diagram pictures a sender AA distributing an $n$-partite GHZ state to herself and (n1)(n-1) other parties. To the left of arrow $(1)$, AA uses a classical protocol to anonymously notify a subset of $m<n$ blue receivers $\{B_{1},B_{2},...,B_{m}\}$ that she will later establish a secret key with them Broadbent and Tapp (2007). To the right of arrow $(1)$, all parties are colored in pink since the shared $n$-partite GHZ state is agnostic of the receivers’ identities. To reach the right of arrow $(2)$, all unnotified parties measure their subsystem in the $X$ basis and broadcast their measurement outcomes in a random order (or through a simultaneous broadcast channel), while the notified parties declare random bits to preserve their anonymity. AA applies a phase flip to her subsystem if the parity of the unnotified parties’ outcomes is odd, yielding the desired $(m+1)$-partite GHZ state.
@@ IX.1 Blind computation
[…]
@@
[…]
@@ IX.2 Verifiable computation
[…]
@@ IX.4 Fully homomorphic encryption
[…]
@@
On the experimental side, a proof-of-principle computation on encrypted quantum data was performed using polarization-encoded single photons and linear optics Fisher et al. (2014). In this scheme, the client is only required to prepare and send single qubits, encrypted through the quantum equivalent of a one-time pad (Section VIII.1). Since client and server agree on the computation beforehand, the overhead in terms of auxiliary qubits is proportional to the number of non-Clifford gates in the computation. Single-qubit Clifford gates are naturally performed using a set of waveplates, while the CNOTs require two-photon interference at a partially polarizing beamsplitter (giving a success probability of $1/9$) and the non-Clifford gates are actively implemented using various Pockels cells to prepare the auxiliary qubit. While a full practical security analysis is not provided, the leakage of information due to imperfect quantum encryption on the client’s side and multiphoton emission from the SPDC source is quantified. Two years later, a similar protocol was implemented in the continuous-variable setting at telecom wavelengths, showing that security can be achieved with up to 1010km of lossy channel between the client and the server Marshall et al. (2016).
@@ X.4 Loss tolerance
X.5 Device independence
While some beyond-QKD primitives have been studied in the measurement-device-independent setting Fu et al. (2015); Li et al. (2023b); Puthoor et al. (2016); Bozzio et al. (2019); Yin et al. (2017); Roberts et al. (2017), the question of full device-independence (DI) Vazirani and Vidick (2014), in which neither quantum sources nor measurement devices can be trusted, remains fairly open. For mistrustful cryptography, DI implies that remote parties must certify the presence of non-locality while displaying competing interests, which differs from the better-studied QKD scenario Acín et al. (2007); Pironio et al. (2009); Woodhead et al. (2021). Although most known mistrustful DI schemes rely on GHZ correlations instead of Bell pairs Silman et al. (2011); Arora et al. (2024), some protocols based on CHSH were proposed for quantum bit commitment and strong coin flipping Aharon et al. (2016). For quantum oblivious transfer, a DI version of the XOR protocol was proposed in Kundu et al. (2022), and security in the bounded storage model was studied under extra post-quantum hardness assumptions Broadbent and Yuen (2023).
Besides raising the question of exactly which quantum-cryptographic primitives admit a DI version and under which assumptions, most existing works are proposed in the ideal setting, where parties do not have to calibrate imperfection parameters related to finite multiphoton emission probability, channel losses, and valid sets of detection time-tags. Thus, in order to develop experimental milestones such as those achieved in DI-QKD Zhang et al. (2022b); Nadlinger et al. (2022), security bounds on the tolerable noise and losses of beyond-QKD primitives should be derived.